OpenStudy (mattduhhelper):

Please show 32/1241 as a Mixed number please.

rebeccaxhawaii (rebeccaxhawaii):

that fraction cant be a mix number it isnt an improper fraction

OpenStudy (mattduhhelper):

Actually it was a originally shown as a division problem tbh

rebeccaxhawaii (rebeccaxhawaii):

so do you mean 1241/32 ?

OpenStudy (mattduhhelper):

Yes!

OpenStudy (imsleeziboii):

38 and 7/10

rebeccaxhawaii (rebeccaxhawaii):

okay so 1241/32 = 38.78125 32 x 38 = 1216 1241-1216= 25 so \[38 \frac{ 25 }{ 32 }\]

OpenStudy (mattduhhelper):

Ok! Thank you <3

OpenStudy (imsleeziboii):

ayee

OpenStudy (mattduhhelper):

?

OpenStudy (imsleeziboii):

:(

OpenStudy (mattduhhelper):

O, well you see i put hers as best response because she told me the equation. No hard feelings! :D

rebeccaxhawaii (rebeccaxhawaii):

no direct answers you need to explain how you got it

OpenStudy (imsleeziboii):

its simple. You divide it

rebeccaxhawaii (rebeccaxhawaii):

theres more to it than just that show how you got the remainder dont expect them to know it